(1 point) Starting salaries of 115 college graduates who have taken a statistics course have a mean of $44,215. Suppose the distribution of this population is approximately normal and has a standard deviation of $10,288. Using a 93% confidence level, find both of the following: (NOTE: Do not use commas or dollar signs in your answers.) (a) The margin of error: equation editor Equation Editor (b) The confidence interval for the mean μ μ
Answer #1

sample mean, xbar = 44215
sample standard deviation, σ = 10288
sample size, n = 115

Given CI level is 93%, hence α = 1 - 0.93 = 0.07
α/2 = 0.07/2 = 0.035, Zc = Z(α/2) = 1.81

a)
ME = zc * σ/sqrt(n)
ME = 1.81 * 10288/sqrt(115)
ME = 1736.44

b)
CI = (xbar - Zc * s/sqrt(n) , xbar + Zc * s/sqrt(n))
CI = (44215 - 1.81 * 10288/sqrt(115) , 44215 + 1.81 * 10288/sqrt(115))
CI = (42478.56 , 45951.44)